The base of a triangle is 3 times its height the area is numerically 6 times its height. Find the height base and area of the triangle.


Answer by ptaylor(2198)


Let x=height of triangle
Then 3x=base of triangle
Area(A)=(1/2)Base(B) times Height(H) or A=(1/2)BH
So our equation to solve is
6x=(1/2)(x)(3x) simplifying
6x=(3x^2)/2 multiply both sides by 2 to get rid of the fraction
12x=3x^2 divide both sides by 3x (x is not equal to zero)
x=4 units-----------------height of triangle
3x=3*4=12 units ---------------base of triangle
Area=(1/2)(4)(12)= 24 sq units
CK
Area is 6 times height
24=6*4
24=24
Base is 3 times height
12=3*4
12=12

Hope this helps---ptaylor